# Information Shapes Koopman Representation

### v1

@@ 1 Introduction
Motivation. This linearization principle has recently attracted significant attention in the deep learning community, as it enables complex nonlinear dynamics to be modeled and predicted using linear representations. However, integrating this framework into deep architectures poses a fundamental challenge: the Koopman operator is inherently infinite-dimensional, necessitating the identification or learning of a suitable finite-dimensional subspace for practical implementation. Deep learning models, most notably variational autoencoders (VAEs), have been employed to approximate such subspaces in a purely data-driven manner (Otto & Rowley, 2019; Pan & Duraisamy, 2020; Liu et al., 2023). Yet in practice, the resulting representations often suffer from instability, mode collapse or fail to produce reliable dynamics. To address these challenges, some studies incorporate domain-specific priors—such as symmetry, conservation laws, dissipation, or ergodicity (Vaidya & Mehta, 2008; Weissenbacher et al., 2022; Azencot et al., 2020; Cheng et al., 2025)—into the Koopman representation. While effective in restricted settings, such approaches lack general principles for guiding Koopman representation. This calls for a more general and principled approach to constructing finite-dimensional representations, one that balances simplicity, in the form of latent linearity, with sufficient expressiveness (more literature review in Appendix C).
Information Bottleneck View. A natural way to achieve the tradeoff between simplicity and expressiveness is through the lens of the Information Bottleneck (IB). The classic IB framework formalizes the idea that a good representation should compress the input as much as possible while preserving information relevant to a downstream task (Tishby et al., 2000; Tishby & Zaslavsky, 2015). In the context of representation learning (see Table 1), this typically means finding a latent variable $z$ that minimizes $\text{Complexity}(x,z)$ [^1] By Complexity(x,z)(x,z) we mean the mutual information $I(x;z)$ in the IB framework, quantifying how much of the input information $x$ is retained in $z$. Relevance(z,y)(z,y) denotes $I(z;y)$, the information $z$ carries about $y$. from input $x$, while retaining expressiveness by improving Relevance(z,y)(z,y) (Vera et al., 2018). Instead of a static latent representation, the goal of Koopman representation is to predict the future state $x_{n}$ given the current state $x_{n-1}$ via a latent variable $z_{n-1}$. This gives rise to a dynamical information bottleneck formulation: we aim to learn a Koopman representation $z_{n-1}$ with maximal linear predictability of future state $x_{n}$, while remaining as compact as possible.
@@
Why Is Finding a Good Koopman Representation Challenging? Learning Koopman representation imposes stricter constraints than conventional latent representation models (see Table 1). In VAE or $\beta$-VAE (Kingma et al., 2013; Burgess et al., 2018), the focus is on reconstructing the input $x$ or sampling from its distribution, which only requires the latent representation $z$ to contain enough information about $y$. However, in Koopman learning, the latent space needs to support linear forward from $z_{n}$ to $z_{n+1}$ in some finite-dimensional spaces. This constraint implies that the latent representation must not only capture information about the current state but also conform to a linear predictive structure (structural consistency), which imposes a stronger restriction. Prior work has shown that simply increasing the dimensions of the latent space does not necessarily improve performance (Li et al., 2020; Brunton et al., 2021), underscoring the importance of maintaining temporal coherence, i.e., ensuring that latent trajectories evolve consistently over time to prevent instability and error accumulation. Moreover, predictive sufficiency requires that the latent representation retains enough Koopman modes to faithfully reconstruct the system’s future trajectories, such that multi-step prediction accuracy is preserved (Wang et al., 2022; 2025). Unlike standard VAEs and their variants, which emphasize flexible latent representations to support reconstruction, Koopman models demand dynamically consistent latent representations: small deviations can propagate and amplify over time. In summary, while conventional representation learning emphasizes disentanglement and reconstruction, Koopman representation learning requires three key properties: temporal coherence, predictive sufficiency, and structural consistency. The IB framework provides a meta view to navigate these trade-offs. It enables us to ask the central question:
@@ 3.2 Information Components in Koopman Representation
The latent mutual information quantifies the magnitude of error, but not the nature of the information lost in the Koopman representation. To sharpen the insight from Propositions 3.1 and 3.1, we consider the aggregated quantity $I(z_{t};x_{t})$, which measures the total information available to the decoder $p(x_{t}|z_{t})$. Our focus is on how much of this information can be stably propagated from past latent variables $z_{t-n}$.
@@
[Effective Dimension and Anti-Collapse] Low effective dimension (see Proposition 1) in Koopman representation indicates information collapse to few dominant modes and limits the model’s ability to represent rich modes. Penalizing the von Neumann entropy $S(\frac{\mathcal{C}}{\mathrm{tr}(\mathcal{C})})$ encourages more expressive and spectrally diverse representations.
Connecting to Proposition 1, Appendix F.6 contains a detailed proof via a water-filling and Lagrangian view. The normalized operator $\frac{\mathcal{C}}{\mathrm{tr}(\mathcal{C})}$ can be regarded as a density matrix in Hilbert space, and the effective dimension can be measured as $\exp(S)$ (see Definition E.2). When penalized with large the von Neumann entropy, the water-filling solution attains a non-zero allocation across all modes, preventing variance from collapsing entirely onto a few dominant directions (cf. equation 46). This ensures a positive distribution of spectral weights across all modes, thereby avoiding degenerate spectra and increasing the effective dimension of the latent space $Z$ (see Figure 1(c)).
@@ 3.3 Information-Theoretic Formulation for Practical Implementation
In VAE structure (shown in Figure 1(a)) , each component of the loss plays a distinct role in balancing the information-theoretic objectives: (1) The mutual information $I(z_{n};\mathcal{P}_{n})$ captures temporal coherence by linking $z_{n}$ to its temporal neighborhood $\mathcal{P}_{n}=\{z_{n\pm i}\mid 1\leq i\leq k\}$, which includes immediate past and future latent states; in practice, this can be computed either via the closed form in equation 7 for low-dimensional latents, or approximated by InfoNCE (Wu et al., 2020) for high-dimensional settings. (2) The term $\mathbb{E}_{p_{\theta}(z_{n}|x_{n})}[\log q_{\psi}(z_{n}|z_{n-1})]-H_{p_{\theta}}(z_{n}|x_{n})$ serves as an equivalent representation of the conditional mutual information $I(z_{t};x_{t}|z_{t-1})$, with linear Gaussian transition $q_{\psi}(z_{n}|z_{n-1})=\mathcal{N}(z_{n}|\mathcal{K}_{\psi}z_{n-1},\Sigma_{\psi})$ and entropy of encoder $H_{p_{\theta}}(z_{n}|x_{n})$ (see Appendix 1). Minimizing this KL not only encourages the latent representation to capture information from the state $x_{n}$, but also compresses fast-dissipating and residual components, ensuring that the representation remains expressive yet simple. Here, $\mathbb{E}_{p_{\theta}(z_{n}|x_{n})}[\log q_{\psi}(z_{n}|z_{n-1})]$ enforces structural consistency in latent space. (3) The term $\log p_{\omega}(x_{n}|z_{n})$ is the decoder loss from predicted latent variable $z_{n}$. (4) von Neumann entropy term $S\!\left(\tfrac{\mathcal{C}}{\mathrm{tr}(\mathcal{C})}\right)$ is computed from the normalized covariance matrix $\mathcal{C}=\tfrac{1}{B}\sum_{i=1}^{B}(z_{i}-\bar{z})(z_{i}-\bar{z})^{\top}$ of the latent codes within a minibatch of size $B$. This promotes spectral diversity and guards against mode collapse, ensuring that the learned Koopman representation retains predictive sufficiency. (5) $\mathcal{L}_{\text{ELBO}}$ is the Evidence Lower Bound (ELBO) for training stability and reconstruction (see more analysis and implementation details in Appendix 1). For AE structure, $\mathbb{E}_{p_{\theta}(z_{n}|x_{n})}[\log q_{\psi}(z_{n}|z_{n-1})]$ degenerates into a $L^{2}$ loss enforcing the structural consistency $\|z_{n+1}-\mathcal{K}_{\psi}z_{n}\|^{2}$, and ELBO becomes AE reconstruction term (see 2).
@@ 4 Experiments
Tasks. We evaluate our approach across three types of dynamical data: (1) Physical simulations, including Lorenz 63, Kármán vortex street, and Dam flow, which test the ability to capture nonlinear high-dimensional physical dynamics; (2) Visual-input control, including image-based Planar, Pendulum, Cartpole, and 3-Link manipulator, which evaluate the ability to extract latent dynamics from high-dimensional visual inputs while controllable in latent spaces; and (3) Graph-structured dynamics prediction, including Rope and Soft Robotics, which tests generalized abilities of latent representation on dynamics with graph structures (see experimental details in Appendix G.3).
Metrics. We assess performance on both forecasting and control. For forecasting, we report (i) normalized root mean square error (NRMSE) for short- and long-term predictions (for physical simulation and graphs-structured dynamics), (ii) physical consistency metrics based on spectral distribution errors (SDEs) based on $1000$-step sequences, (iii) distributions of state measured by the Kullback–Leibler divergence (KLD), and (iv) structural similarity index measure (SSIM) for physical simulations. (v) the quality of low-dimensional manifold construction from high-dimensional visual inputs. For control, we measure the success rate of latent-space control of visual inputs following the setting in (Levine et al., 2020).
@@
(3) How does explicit information-theoretic regularization sufficiently capture essential dynamics, compared with VAEs and Koopman autoencoders? (Proposition 1 and 1) As shown in the reconstructed manifolds of Figure 3, our method produces a latent manifold that aligns most closely with the ground truth. For E2C, which is directly built on a VAE architecture, the latent geometry is heavily distorted (the loss of coherence). The manifold learned by KAE collapses into a nearly one-dimensional structure, reflecting the lack of effective dimensions in its latent space. PCC, a modified VAE-based method designed to improve manifold construction, demonstrates partial improvement but still exhibits a gap compared with our approach. By preserving both effective dimensionality and temporal coherence, our Koopman representation achieves the best average control performance in both noiseless and noisy environments (Table 6 and 7 in Appendix G.5.2).
@@
(4) How robust are the findings under noise and longer horizons? (Proposition 3.1 and 3.1) Our method remains robust under both noisy observations and extended prediction horizons. As shown in Table 2, it maintains stable performance in long-term rollouts and physical statistics. Moreover, our approach supports control under noisy environments, achieving competitive performance. These quantitative results are consistent with our probabilistic propositions.
(5) To what extent can our Lagrangian formulation be generalized to diverse architectures and adapted to support downstream tasks? (Proposition 3.1-1) Our formulation demonstrates broad applicability: it consistently improves performance across physical simulations (see Table 2), visual perception tasks for manifold construction and control (see Figure 3, Tables 6 and 7 in Appendix G.5.2), and graph-structured dynamics prediction (see Figure 4). These gains indicate that the proposed Lagrangian principle is architecture-agnostic and can be readily incorporated into different settings to enhance both predictive accuracy and task effectiveness (more results are referred to Appendix G.5).
@@ Appendix C More Literature Review Related to Koopman Representation
The Koopman operator was originally introduced by Koopman and von Neumann as a linear embedding of Hamiltonian dynamical systems (Koopman, 1931; Koopman & Neumann, 1932). However, its infinite-dimensional nature makes it difficult to identify suitable handcrafted basis functions using conventional methods (Brunton et al., 2021). To address this, kernel techniques from functional analysis have been employed as bases for learning the Koopman operator (Das & Giannakis, 2020; Das et al., 2021; Kostic et al., 2022; Bevanda et al., 2025). Owing to the well-posed properties of kernel functions in reproducing kernel Hilbert spaces (e.g., linearity, existence, and convergence guarantees), the Koopman operator can be directly approximated via (extended) dynamic mode decomposition (DMD or EDMD) (Williams et al., 2015; Takeishi et al., 2017; Arbabi & Mezic, 2017). Despite these theoretical advantages, fixed kernel functions are often too restrictive to capture a general function space (Berlinet & Thomas-Agnan, 2011; Alpay, 2012).
In contrast, deep learning frameworks provide a more flexible alternative: leveraging the universal approximation property of neural networks (Baker & Patil, 1998; Kidger & Lyons, 2020), they allow learning a general Koopman representation without relying on predefined kernels. Following this principle, (variational) autoencoder (AE/VAE) architectures have been widely adopted to extract features spanning the Koopman subspace (Liu et al., 2023; Wu et al., 2025). The resulting latent representations are flexible and support downstream tasks such as prediction and control (Li et al., 2020; Mauroy et al., 2020; Korda & Mezić, 2020; Weissenbacher et al., 2022). However, these representations are typically learned in a purely self-supervised manner, lacking explicit grounding in dynamical systems theory. To improve their reliability, recent studies incorporate domain-specific priors—such as symmetry, conservation laws, dissipation, or ergodicity—into the Koopman representation (Vaidya & Mehta, 2008; Weissenbacher et al., 2022; Azencot et al., 2020; Cheng et al., 2025). While existing approaches are effective for specific dynamical systems, a formal theoretical foundation for guiding the learning of Koopman representations remains insufficient. In this work, we investigate how general information-theoretic principles can be employed to fill this gap.

### v2

@@ 1 Introduction
Motivation. This linearization principle has recently attracted significant attention in the deep learning community, as it enables complex nonlinear dynamics to be modeled and predicted using linear representations. However, integrating this framework into deep architectures poses a fundamental challenge: the Koopman operator is inherently infinite-dimensional, necessitating the identification or learning of a suitable finite-dimensional subspace for practical implementation. Deep learning models, most notably variational autoencoders (VAEs), have been employed to approximate such subspaces in a purely data-driven manner (Otto and Rowley, 2019; Pan and Duraisamy, 2020; Liu et al., 2023). Yet in practice, the resulting representations often suffer from instability, mode collapse or fail to produce reliable dynamics. To address these challenges, some studies incorporate domain-specific priors—such as symmetry, conservation laws, dissipation, or ergodicity (Vaidya and Mehta, 2008; Weissenbacher et al., 2022; Azencot et al., 2020; Cheng et al., 2025)—into the Koopman representation. While effective in restricted settings, such approaches lack general principles for guiding Koopman representation. This calls for a more general and principled approach to constructing finite-dimensional representations, one that balances simplicity, in the form of latent linearity, with sufficient expressiveness (more literature review in Appendix C).
Information Bottleneck View. A natural way to achieve the tradeoff between simplicity and expressiveness is through the lens of the Information Bottleneck (IB). The classic IB framework formalizes the idea that a good representation should compress the input as much as possible while preserving information relevant to a downstream task (Tishby et al., 2000; Tishby and Zaslavsky, 2015). In the context of representation learning (see Table 1), this typically means finding a latent variable $z$ that minimizes $\text{Complexity}(x,z)$ [^1] By Complexity(x,z)(x,z) we mean the mutual information $I(x;z)$ in the IB framework, quantifying how much of the input information $x$ is retained in $z$. Relevance(z,y)(z,y) denotes $I(z;y)$, the information $z$ carries about $y$. from input $x$, while retaining expressiveness by improving Relevance(z,y)(z,y) (Vera et al., 2018). Instead of a static latent representation, the goal of Koopman representation is to predict the future state $x_{n}$ given the current state $x_{n-1}$ via a latent variable $z_{n-1}$. This gives rise to a dynamical information bottleneck formulation: we aim to learn a Koopman representation $z_{n-1}$ with maximal linear predictability of future state $x_{n}$, while remaining as compact as possible.
@@
Why Is Finding a Good Koopman Representation Challenging? Learning Koopman representation imposes stricter constraints than conventional latent representation models (see Table 1). In VAE or $\beta$-VAE (Kingma et al., 2013; Burgess et al., 2018), the focus is on reconstructing the input $x$ or sampling from its distribution, which only requires the latent representation $z$ to contain enough information about $y$. However, in Koopman learning, the latent space needs to support linear forward from $z_{n}$ to $z_{n+1}$ in some finite-dimensional spaces. This constraint implies that the latent representation must not only capture information about the current state but also conform to a linear predictive structure (structural consistency) (Mardt et al., 2018; Kostic et al., 2023a; b; 2024), which imposes a stronger restriction. Prior work has shown that simply increasing the dimensions of the latent space does not necessarily improve performance (Li et al., 2020; Brunton et al., 2021), underscoring the importance of maintaining temporal coherence, i.e., ensuring that latent trajectories evolve consistently over time to prevent instability and error accumulation. Moreover, predictive sufficiency requires that the latent representation retains enough Koopman modes to faithfully reconstruct the system’s future trajectories, such that multi-step prediction accuracy is preserved (Wang et al., 2022; 2025). Unlike standard VAEs and their variants, which emphasize flexible latent representations to support reconstruction, Koopman models demand dynamically consistent latent representations: small deviations can propagate and amplify over time. In summary, while conventional representation learning emphasizes disentanglement and reconstruction, Koopman representation learning requires three key properties: temporal coherence, predictive sufficiency, and structural consistency. The IB framework provides a meta view to navigate these trade-offs. It enables us to ask the central question:
@@ 3.2 Information Components in Koopman Representation
The latent mutual information quantifies the magnitude of error, but does not uncover how this loss relates to Koopman spectral properties. To sharpen the insight from Propositions 3.1 and 3.1, we consider the aggregated quantity $I(z_{t};x_{t})$, which measures the total information available to the decoder $p(x_{t}|z_{t})$. Our focus is on how much of this information can be stably propagated from past latent variables $z_{t-n}$.
@@
[Effective Dimension and Anti-Collapse] Low effective dimension (see Proposition 1) in Koopman representation indicates information collapse to few dominant modes and limits the model’s ability to represent rich modes. Penalizing the von Neumann entropy $S(\frac{\mathcal{C}}{\mathrm{tr}(\mathcal{C})})$ encourages more expressive and spectrally diverse representations. Connecting to Proposition 1, Appendix F.6 contains a detailed proof via a water-filling and Lagrangian view. The normalized operator $\frac{\mathcal{C}}{\mathrm{tr}(\mathcal{C})}$ can be regarded as a density matrix in Hilbert space, and the effective dimension can be measured as $\exp(S)$ (see Definition E.2). When penalized with large the von Neumann entropy, the water-filling solution attains a non-zero allocation across all modes, preventing variance from collapsing entirely onto a few dominant directions (cf. equation 46). This ensures a positive distribution of spectral weights across all modes, thereby avoiding degenerate spectra and increasing the effective dimension of the latent space $Z$ (see Figure 1(c)).
@@ 3.3 Information-Theoretic Formulation for Practical Implementation
In VAE structure (shown in Figure 1(a)) , each component of the loss plays a distinct role in balancing the information-theoretic objectives: (1) The mutual information $I(z_{n};\mathcal{P}_{n})$ captures temporal coherence by linking $z_{n}$ to its temporal neighborhood $\mathcal{P}_{n}=\{z_{n\pm i}\mid 1\leq i\leq k\}$, which includes immediate past and future latent states; in practice, this can be computed either via the closed form in equation 7 for low-dimensional latents, or approximated by InfoNCE (Wu et al., 2020) for high-dimensional settings. (2) The term $\mathbb{E}_{p_{\theta}(z_{n}|x_{n})}[\log q_{\psi}(z_{n}|z_{n-1})]-H_{p_{\theta}}(z_{n}|x_{n})$ serves as an equivalent representation of the conditional mutual information $I(z_{t};x_{t}|z_{t-1})$, with linear Gaussian transition $q_{\psi}(z_{n}|z_{n-1})=\mathcal{N}(z_{n}|\mathcal{K}_{\psi}z_{n-1},\Sigma_{\psi})$ and entropy of encoder $H_{p_{\theta}}(z_{n}|x_{n})$ (see Appendix G.1). Minimizing this KL not only encourages the latent representation to capture information from the state $x_{n}$, but also compresses fast-dissipating and residual components, ensuring that the representation remains expressive yet simple. Here, $\mathbb{E}_{p_{\theta}(z_{n}|x_{n})}[\log q_{\psi}(z_{n}|z_{n-1})]$ enforces structural consistency in latent space. (3) The term $\log p_{\omega}(x_{n}|z_{n})$ is the decoder loss from predicted latent variable $z_{n}$. (4) von Neumann entropy term $S\!\left(\tfrac{\mathcal{C}}{\mathrm{tr}(\mathcal{C})}\right)$ is computed from the normalized covariance matrix $\mathcal{C}=\tfrac{1}{B}\sum_{i=1}^{B}(z_{i}-\bar{z})(z_{i}-\bar{z})^{\top}$ of the latent codes within a minibatch of size $B$. This promotes spectral diversity and guards against mode collapse, ensuring that the learned Koopman representation retains predictive sufficiency. (5) $\mathcal{L}_{\text{ELBO}}$ is the Evidence Lower Bound (ELBO) for training stability and reconstruction (see more analysis and implementation details in Appendix G.1). For AE structure, $\mathbb{E}_{p_{\theta}(z_{n}|x_{n})}[\log q_{\psi}(z_{n}|z_{n-1})]$ degenerates into a $L^{2}$ loss enforcing the structural consistency $\|z_{n+1}-\mathcal{K}_{\psi}z_{n}\|^{2}$, and ELBO becomes AE reconstruction term (see G.2).
@@ 4 Experiments
Tasks. We evaluate our approach across three types of dynamical data: (1) Physical simulations, including Lorenz 63, Kármán vortex street, Dam flow, and weather forecasting task (ERA5), which test the ability to capture nonlinear, stochastic and high-dimensional physical dynamics; (2) Visual-input control, including image-based Planar, Pendulum, Cartpole, and 3-Link manipulator, which evaluate the ability to extract latent dynamics from high-dimensional visual inputs while controllable in latent spaces; and (3) Graph-structured dynamics prediction, including Rope and Soft Robotics, which tests generalized abilities of latent representation on dynamics with graph structures (see experimental details in Appendix G.3).
Metrics. We assess performance on both forecasting and control. For forecasting, we report (i) normalized root mean square error (NRMSE) for short- and long-term predictions (for physical simulation and graphs-structured dynamics), (ii) physical consistency metrics based on spectral distribution errors based on $1000$-step sequences (SDEs), (iii) distributions of state measured by the Kullback–Leibler divergence (KLD), and (iv) structural similarity index (SSIM) for physical simulations. (v) the quality of low-dimensional manifold construction from high-dimensional visual inputs. For control, we measure the success rate of latent-space control of visual inputs following the setting in (Levine et al., 2020).
@@
ERA5 Weather (channel avg) 5-NRMSE 0.055 0.058 0.049 0.028
10-NRMSE 0.063 0.068 0.060 0.035
50-NRMSE 0.118 0.074 0.079 0.068
5-SSIM 0.666 0.664 0.697 0.867
10-SSIM 0.619 0.606 0.635 0.808
50-SSIM 0.481 0.707 0.695 0.781
@@
(3) How does explicit information-theoretic regularization sufficiently capture essential dynamics, compared with VAEs and Koopman autoencoders? (Proposition 1 and 1) As shown in the reconstructed manifolds of Figure 3, our method produces a latent manifold that aligns most closely with the ground truth. For E2C, which is directly built on a VAE architecture, the latent geometry is heavily distorted (the loss of coherence). The manifold learned by KAE collapses into a nearly one-dimensional structure, reflecting the lack of effective dimensions in its latent space. PCC, a modified VAE-based method designed to improve manifold construction, demonstrates partial improvement but still exhibits a gap compared with our approach. By preserving both effective dimensionality and temporal coherence, our Koopman representation achieves the best average control performance in both noiseless and noisy environments (Table 8 and 9 in Appendix G.5.2).
@@
Figure 4: Comparison of continuous predictions for the global humidity starting from 2018010100:002018-01-01-00:$0$ to 2018010808:002018-01-08-08:$0$. Error maps in the lower panels demonstrate that, compared with other models, showing with more stable and accurate results of our model (see more demonstration in Appendix G.5.1).
(4) How robust are the findings under noise, extended prediction horizons, and large-scale settings? (Proposition 3.1 and 3.1) Our method remains robust under both noisy observations and extended prediction horizons. As shown in Table 2 and Figure 4, it maintains stable performance in long-term rollouts and physical statistics in large scale weather forecasting. Moreover, our approach supports control under noisy environments, achieving competitive performance. These quantitative results are consistent with our probabilistic propositions.
(5) To what extent can our Lagrangian formulation be generalized to diverse architectures and adapted to support downstream tasks? (Proposition 3.1-1) Our formulation demonstrates broad applicability: it consistently improves performance across physical simulations (see Table 2), visual perception tasks for manifold construction and control (see Figure 3, Tables 8 and 9 in Appendix G.5.2), and graph-structured dynamics prediction (see Figure 5). These gains indicate that the proposed Lagrangian principle is architecture-agnostic and can be readily incorporated into different settings to enhance both predictive accuracy and task effectiveness (more results are referred to Appendix G.5).
@@ 5 Conclusion
ACKNOWLEDGMENTS
Zhuo Sun is supported by Fundamental Research Funds for the Central Universities 2025110590 of Shanghai University of Finance and Economics.
Ethics Statement and Reproducibility Statement
This work raises no specific ethical concerns beyond standard practices in machine learning research. All methods, datasets, and hyperparameters are described in detail, and core code is released in supplementary materials.
@@ Appendix C More Literature Review Related to Koopman Representation
The Koopman operator was originally introduced by Koopman and von Neumann as a linear embedding of Hamiltonian dynamical systems (Koopman, 1931; Koopman and Neumann, 1932). However, its infinite-dimensional nature makes it difficult to identify suitable handcrafted basis functions using conventional methods (Brunton et al., 2021). To address this, kernel techniques from functional analysis have been employed as bases for learning the Koopman operator (Das and Giannakis, 2020; Das et al., 2021; Kostic et al., 2022; Bevanda et al., 2025). Owing to the well-posed properties of kernel functions in reproducing kernel Hilbert spaces (e.g., linearity, existence, and convergence guarantees), the Koopman operator can be directly approximated via (extended) dynamic mode decomposition (DMD or EDMD) (Williams et al., 2015; Takeishi et al., 2017; Arbabi and Mezic, 2017; Xu et al., 2025a). Despite these theoretical advantages, fixed kernel functions are often too restrictive to capture a general function space (Berlinet and Thomas-Agnan, 2011; Alpay, 2012).
In contrast, deep learning frameworks provide a more flexible alternative: leveraging the universal approximation property of neural networks (Baker and Patil, 1998; Kidger and Lyons, 2020), they allow learning a general Koopman representation without relying on predefined kernels. Following this principle, (variational) autoencoder (AE/VAE) architectures have been widely adopted to extract features spanning the Koopman subspace (Liu et al., 2023; Wu et al., 2025; Xu et al., 2025b; c). The resulting latent representations are flexible and support downstream tasks such as prediction and control (Li et al., 2020; Mauroy et al., 2020; Korda and Mezić, 2020; Weissenbacher et al., 2022). However, these representations are typically learned in a purely self-supervised manner, lacking explicit grounding in dynamical systems theory. To improve their reliability, recent studies incorporate domain-specific priors—such as symmetry, conservation laws, dissipation, or ergodicity—into the Koopman representation (Vaidya and Mehta, 2008; Weissenbacher et al., 2022; Azencot et al., 2020; Cheng et al., 2025). Within the VAE setting, recent studies (Federici et al., 2023) have started to link Markovian dynamics and information theory, demonstrating that time-lagged tricks can exploit mutual information to obtain better latent representations. While existing approaches are effective for specific dynamical systems, a formal theoretical foundation for guiding the learning of Koopman representations remains insufficient. In this work, we investigate how general information-theoretic principles can be employed to fill this gap.